Find the inverse function of f(x)= 4-x^2, x=> 0
Answer by amarjeeth123(574)

Let y= 4-x^2
Replacing x with y and y with x we get,
x=4-y^2
y^2=4-x
y=√(4-x)
F(x)=√(4-x) is the inverse function of the given function.